OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form.

OpenStudy (hba):

x/a+y/b=1

OpenStudy (anonymous):

y = 7x – 10 y = 7x + 10 y = 7x + 18 y = 7x – 18 are the anwsers i had got

OpenStudy (anonymous):

which one is correct

OpenStudy (hba):

@SmoothMath please help him out

OpenStudy (anonymous):

You got 4 different answers but you're not sure which is correct? How did you get 4 different answers?

OpenStudy (anonymous):

nvm got the anwser

OpenStudy (anonymous):

Oh yeah?